The following data set is incomplete.
27, 29, 27, 25, 24, 27, 25, 29

Four identical values, all below 23 are missing.
(i) Work out the median.
(ii) Make a comment about the mode.
(iii) The mean of all 12 values is known to be 24.75.
What are the missing values ?

Answer by ewatrrr(24785)

 
Hi
x, x, x, x, 24, 25, 25, 27, 27, 27, 29, 29
median = (25+25)/2 = 25
Mode = 4 (unknown identical items)
(x + x + x + x + 24 + 25 + 25 + 27 + 27 + 27 + 29 + 29)/12 = 24.75
4x = 12*24.75 - (24 + 25 + 25 + 27 + 27 + 27 + 29 + 29)
x = 3*24.75 - (24 + 25 + 25 + 27 + 27 + 27 + 29 + 29)